OpenStudy (anonymous):

You put $200 in an account with 6% APR for ten years. Find the total savings if interest is compounded a) annually, b)monthly, c)daily, and d) continously

OpenStudy (anonymous):

A=P(1+r/n)^(n*t) P for priciple, t for time , r for rate , a) A= 200(1+(0.06/2))^(2*10) A= $361.22 b) A= 200(1+0.06/12)^(12*10) A= $363.9 c) A= 200(1+0.06/365)^(365*10) A = $364.4 \[d) 200*e^{(0.06*10)} = \$364,42\]